OpenStudy (anonymous):

Please Help me with this question: Which is evivalent to 17.5mm A. 0.175cm B. 1.75cm C.175cm D.1750cm

OpenStudy (anonymous):

1.75cm

OpenStudy (anonymous):

B

OpenStudy (anonymous):

17.5mm=17.5*10^-3m=0.0175m=1.75*10^-2m=1.75cm milli=10^-3 cent=10^-2 does that help your understanding?

OpenStudy (anonymous):

thank you everyone :)